# Harnessing the Unseen: The Hidden Influence of Intrinsic Knowledge in Long-Context Language Models

### v1

@@ 4.2 Experiment
Results on HotpotQA-context subset and HotpotQA-internal subset are shown in Figure 3. STRING significantly improves performance on the HotpotQA-context subset, where the model needs to rely on the provided context to derive the reference answer. The improved performance implies that STRING is improving extrinsic retrieval ability on LCMs which supports our hypothesis above. However, on the HotpotQA-internal subset where the context is related but not useful, we find STRING underperforms RoPE which matches our finding on I-WhoQA-irrelevant. In conclusion, the enhanced extrinsic retrieval ability of STRING interferes with the expression of the model’s intrinsic knowledge, resulting in worse performance compared to the vanilla models. This trend becomes even more apparent on longer context length since STRING’s boosted extrinsic retrieval ability can be more prominent with longer contexts as shown on the Llama-3.1-8B-Instruct HotPotQA-context subset.
5 Hybrid Needle-in-a-Haystack Framework
So far, we have confirmed that intrinsic knowledge plays a crucial role in the generation process of LCMs, but the intrinsic retrieval ability does not improve in parallel with the extrinsic retrieval ability. In this section, we want to examine LCMs’ ability to integrate the intrinsic and extrinsic retrieval ability during long-context generation. Previous benchmarks (Hsieh et al., 2024; An et al., 2023; Zhang et al., 2024) have primarily focused on the effective utilization of external context and overlooked assessing the impact of LCMs’ intrinsic knowledge. To address this gap, we designed a simple yet effective Hybrid Needle-in-a-Haystack (NIAH) test aimed at simultaneously testing intrinsic and extrinsic retrieval ability, providing a more comprehensive assessment of how these two capabilities interact, complement, or potentially interfere with one another, ultimately affecting the quality of generation.
5.1 Dataset Creation
NIAH test uses irrelevant content as a haystack and inserts the answer from a question-answer pair into the haystack to evaluate LCMs’ extrinsic retrieval ability under different context lengths and insertion depths. Unlike the original NIAH, our Hybrid NIAH test simultaneously tests both retrieval abilities by integrating scenarios where the model needs to first utilize its intrinsic knowledge and then perform extrinsic retrieval on the haystack. A comparison between NIAH and Hybrid NIAH is illustrated in Figure 4. The Hybrid NIAH introduces an extra step designed to trigger the intrinsic retrieval ability, requiring the model to recall the relevant target from intrinsic knowledge. The model must use this recalled target to perform extrinsic retrieval from the provided context, ultimately producing the final output.
Firstly, to implement the intrinsic retrieval step, we use the I-WhoQA dataset constructed in Section 3. To ensure that the question-answer pairs only require knowledge that exists across all models, we perform an intersection operation between the I-WhoQA datasets obtained from Mistral-7B-Instruct-v0.2 and Llama3.1-8B-Instruct. Secondly, for the extrinsic retrieval step, we adopt a similar approach to the NIAH test and create various random facts to serve as inserted answers to the corresponding questions. For the remaining irrelevant haystack construction, we also follow the NIAH test and use the PaulGrahamEssays dataset  [^1] https://huggingface.co/datasets/sgoel9/paul_graham_essays to construct the target length haystack.
To prevent the model from exploiting potential syntactic patterns present in the question-answer pairs to extract the inserted “needle” from the context, we introduce varying amounts of random facts as interference (from 0 to 3). This additional noise ensures that models understand the task and the context rather than relying on superficial patterns, thereby enhancing the robustness and reliability of the final result. Details of the final knowledge pairs are listed in Appendix A.
Figure 4: Upper) Needle-in-a-Haystack. It involves directly inserting the answer into the haystack and retrieving it. Lower) Hybrid Needle-in-a-Haystack. It requires a two-step process: first, performing intrinsic retrieval to identify the retrieval target based on the model’s intrinsic knowledge, and then retrieving the answer from the haystack.
5.2 Experiment
Experimental Setting
We evaluated models from three different families with various specifications: Mistral, Llama3.1, and Qwen2.5. For all models, we utilized the Instruct versions and employed greedy decoding strategy. We set the maximum generation length to 32 and 64. For each model, we divided the maximum context length into 40 intervals. For each interval, we defined 10 progressively increasing insertion depths and conducted Hybrid NIAH tests accordingly. This setup allows us to systematically evaluate the models’ performance across varying input lengths and insertion depths. To ensure the stability of the results, we use multiple examples and report the average result. Finally, to measure the similarity between the predictions and the references, we adopt the method provided by PyramidKV (Cai et al., 2024) [^2] https://github.com/Zefan-Cai/KVCache-Factory/, which calculates the final result based on exact matches between tokens. The formula is as follows:
$Score=\frac{|P|\cap|A|}{|A|}$
where $P$ and $A$ are tokens from predictions and references respectively. More detail settings are listed in Appendix Table 2.
@@ 5.2 Experiment
The averaged results for the Hybrid NIAH test are presented in Table 1 and we list the detailed results in Appendix B. From these results, we can observe three interesting findings. Firstly, our Hybrid NIAH test is significantly more challenging than the standard NIAH test, especially on the long-context scenarios as shown in the results on Qwen2.5-7B-Instruct-1M in Figure 5. Moreover, by inserting varying amounts of random facts, we observe a gradual decline in model performance, while the NIAH test results remain unaffected by this interference as shown in Table 1. This discrepancy indicates that models tend to rely on superficial patterns within the context to extract answers when there is a single needle, rather than genuinely leveraging their intrinsic knowledge. The insertion of random facts effectively disrupts this pattern-based retrieval process, allowing us to better assess whether a model truly utilizes its intrinsic knowledge rather than merely exploiting contextual shortcuts. Multiple-needle Hybrid NIAH test serves as a more rigorous evaluation of how well models integrate their intrinsic retrieval ability with the long-context generation.
@@
Secondly, models from the Mistral and Llama3.1 families fail to achieve significant improvements on the Hybrid-NIAH test, even when upgraded to larger or more advanced versions. This suggests that these models struggle to effectively utilize their intrinsic knowledge under long-context scenarios. Despite increased model capacity, their performance remains constrained when external context is incomplete or ambiguous. In contrast, the Qwen2.5 family shows substantial improvements as the model scales up. Larger Qwen2.5 models are more capable of effectively leveraging their intrinsic knowledge, providing more accurate and factually consistent results even when the external context is incomplete.
Figure 5: Hybrid NIAH test results. Upper) Qwen2.5-7B-Instruct-1M with generation length 32. Lower) Qwen2.5-72B-Instruct with generation lengths 32 and 64. The number of random facts was set to 0.
Finally, we observe an interesting phenomenon in the Qwen2.5 models: when the generation length is set to 32, both Qwen2.5-14B-Instruct and Qwen2.5-72B-Instruct perform worse that Qwen2.5-7B-Instruct on the single-needle Hybrid NIAH test. However, when the generation length is increased to 64, their performance improves dramatically as shown in Figure 5. Upon inspecting the outputs manually, we find that these models often first explicitly state their inability to retrieve the answer from the context before producing the correct result. Interestingly, when we introduce several random facts, the models can directly answer the corresponding question without hesitation as shown in Appendix Figure 7. We hypothesize that this behavior is closely related to the training process of Qwen2.5 family. Firstly, when only a single fact is present, the model’s chunked attention mechanism (An et al., 2024a) struggle to effectively focus on the inserted needle when generating the first token. As a result, larger Qwen variants initially rely on their logical reasoning to realize that they do not know the answer and start generating a refusal. But as each refusal token is generated, the model can attend to the context again to find the needle and generate the correct answer even through it refused initially. In contrast, the smaller Qwen-2.5-7B-Instruct model starts hallucinating an answer although it’s unsure instead of outright refusing. Secondly, when multiple needles are inserted, the syntactical similarity between the needles help the larger Qwen models to focus on the context and they do not refuse answering initially. This observation suggests that the Qwen2.5 attention mechanism is highly sensitive to pattern similarity, and syntactical structure can help guide the model’s attention to the relevant part from the context. We leave further investigation of this phenomenon for future work.
6 Conclusion
In this paper, we reveals that the intrinsic knowledge within LCMs plays a crucial role in generation, particularly under long-context scenarios. We demonstrate that enhancing extrinsic retrieval ability can inadvertently suppress the model’s ability to utilize its intrinsic knowledge, highlighting a fundamental trade-off between these two retrieval mechanisms. We further propose a Hybrid Needle-in-a-Haystack test, providing a comprehensive framework to evaluate both intrinsic and extrinsic retrieval ability. Our findings show that even large LCMs struggle to combine the two retrieval abilities and cannot fully leverage its intrinsic knowledge during the long-context generation.
Supplementary Material: Appendices
Figure 6: One example on I-WhoQA dataset on Mistral-7B-Instruct-v0.2.
Model Num of Examples Max Context Length Num of Depths Num of Intervals
Mistral-7B-Instruct-v0.2 5 32768 40 10
Mistral-7B-Instruct-v0.3 2 32768 40 10
Llama-3.1-8B-Instruct 5 131072 40 10
Llama-3.1-70B-Instruct 2 131072 40 10
Qwen2.5-7B-Instruct 5 131072 40 10
Qwen2.5-14B-Instruct 2 131072 40 10
Qwen2.5-72B-Instruct 2 131072 40 10
Qwen2.5-7B-Instruct-1M-128k 5 131072 40 10
Qwen2.5-7B-Instruct-1M 2 1010000 40 10
Table 2: NIAH and Hybrid NIAH test configurations for different backend models.
Figure 7: Hybrid NIAH examples from Qwen2.5-72B-Instruct. The generation length was set to 64.
@@ Appendix A Hybrid Needle-in-a-Haystack Knowledge Dataset
All entities used in Hybrid NIAH tests are illustrated in Figure 3, with the bolded entries representing the target entities used for generating the desired answers. To construct the random facts required for our Hybrid NIAH tests, we randomly select entities from all entities other than the current target entity.

### v2

@@ Introduction
Results on HotpotQA-Context subset and HotpotQA-Parametric subset are shown in Figure 2. STRING significantly improves performance on the HotpotQA-Context subset, where the model must rely on the external context to derive the correct answer. This confirmed that STRING is improving extrinsic retrieval ability on LCLMs. However, on the HotpotQA-Parametric subset where the context is related but not useful, STRING underperforms RoPE and this is consistent with our finding on the I-WhoQA-Irrelevant subset. In conclusion, STRING’s enhanced extrinsic retrieval ability interferes with the model’s use of parametric knowledge, leading to degraded performance in scenarios where the context is misleading or insufficient. This trade-off becomes more pronounced as the context length increases since STRING’s boosted extrinsic retrieval ability can be more prominent with longer contexts as shown on the Llama-3.1-8B-Instruct HotPotQA-Context subset.
@@
Hybrid Needle-in-a-Haystack Framework
So far, we have shown that parametric knowledge plays a crucial role in the long-context generation, yet improvements in extrinsic retrieval do not translate into better parametric recall. Moreover, existing benchmarks (Hsieh et al. 2024; An et al. 2023; Zhang et al. 2024) primarily focus on how to utilize external context, overlooking scenarios that require interaction with parametric knowledge. As a result, methods (An et al. 2024b) developed under these benchmarks fail to achieve better performance when parametric recall is essential. To address this gap, we propose a simple yet effective Hybrid Needle-in-a-Haystack (NIAH) test designed to jointly evaluate parametric recall ability and extrinsic retrieval ability. This test enables a more comprehensive assessment of how these two forms of knowledge interact, complement, or interfere with one another, ultimately affecting the quality of model generation.
Dataset Creation
NIAH test uses irrelevant content as a haystack and inserts the answer from a question-answer pair into the haystack to evaluate LCLMs’ extrinsic retrieval ability under different context lengths and insertion depths. Unlike the standard NIAH test, our Hybrid NIAH test simultaneously assesses parametric recall ability and extrinsic retrieval ability by designing scenarios where the model must first utilize its parametric knowledge and then perform extrinsic retrieval on the haystack. A comparison between NIAH test and Hybrid NIAH test is illustrated in Figure 4. The Hybrid NIAH test introduces an additional step designed to trigger the parametric recall ability, requiring the model to recall the relevant target from parametric knowledge. The model relies on this recalled target entity to guide extrinsic retrieval from the external context, ultimately producing the final answer.
Firstly, to implement the parametric recall step, we use the I-WhoQA dataset constructed in Section 1. To ensure that the question-answer pairs only require knowledge that exists across all models, we perform an intersection operation between the I-WhoQA datasets obtained from Mistral-7B-Instruct-v0.2 and Llama3.1-8B-Instruct. Secondly, for the extrinsic retrieval step, we adopt a similar approach to the NIAH test and create various random facts to serve as inserted answers to the corresponding questions. For the remaining irrelevant haystack construction, we follow the NIAH test and use the PaulGrahamEssays dataset to construct the target length haystack.
To prevent the model from exploiting potential syntactic patterns present in the question-answer pairs to extract the inserted “needle” from the context, we introduce varying amounts of random facts as interference (from 0 to 3). This additional noise ensures that models understand the task and the context rather than relying on superficial patterns, thereby enhancing the robustness and reliability of the final result. Details of the final knowledge pairs are listed in the Appendix Section 1.
Experiment
Experimental Setting
We evaluated models from three different families with various specifications: Mistral, Llama3.1, and Qwen2.5. For all models, we utilized the Instruct versions and employed greedy decoding strategy. We set the maximum generation length to 32 and 64. For each model, we divided the maximum context length into 40 intervals. For each interval, we defined 10 progressively increasing insertion depths and conducted Hybrid NIAH tests accordingly. This setup allows us to systematically evaluate the models’ performance across varying context lengths and insertion depths. To ensure the stability of the results, we use multiple examples and report the average result. Finally, to measure the similarity between the predictions and the references, we adopt the method provided by PyramidKV (Cai et al. 2024), which calculates the final result based on exact matches between tokens. The formula is as follows:
$Score=\frac{|P|\cap|A|}{|A|}$
where $P$ and $A$ are tokens from predictions and references respectively. More detail settings are listed in the Appendix Table 2.
Figure 5: Hybrid NIAH test results. Upper) Qwen2.5-7B-Instruct-1M with generation length 32 and various numbers of inserted random facts. Lower) Qwen2.5-72B-Instruct with generation lengths 32 and 64. The number of random facts was set to 0.
@@
The averaged results for the Hybrid NIAH test are presented in Table 1 and we list the detailed results in the Appendix Section 2. From these results, we can observe three interesting findings.
@@
Firstly, with only a simple modification to the question, our Hybrid NIAH test becomes significantly more challenging than the standard NIAH test, as shown in Table 1. This effect is especially evident in long-context scenarios, as demonstrated by the results on Qwen2.5-7B-Instruct-1M in Figure 5. Moreover, by inserting varying amounts of random facts, we observe a clear decline in model performance by up to 25% for Llama-3.1-70B-Instruct, while the NIAH test results remain unaffected by this interference, as shown in Table 1. This discrepancy indicates that models rely on superficial patterns within the context to extract answers when there is a single needle, rather than genuinely leveraging their parametric knowledge. The insertion of random facts effectively disrupts this pattern-based retrieval process, allowing us to better assess whether a model truly utilizes its parametric knowledge rather than merely exploiting contextual shortcuts. Multiple-Needles Hybrid NIAH test serves as a more rigorous evaluation of how well models integrate their parametric recall ability with the long-context generation.
@@
Secondly, models from the Mistral and Llama3.1 families fail to achieve significant improvements on the Hybrid-NIAH test, even when upgraded to larger versions. This suggests that these models struggle to effectively utilize their parametric knowledge under long-context scenarios. Despite increased model capacity, their performance remains constrained when external context is incomplete or ambiguous. In contrast, the Qwen2.5 family shows substantial improvements as model size scales up. Larger Qwen2.5 models are more capable of effectively leveraging their parametric knowledge, providing more accurate and factually consistent results even when the external context is incomplete.
Finally, we observe an interesting phenomenon in the Qwen2.5 models for Single-Needle Hybrid NIAH test: larger model variants initially claim to not see the inserted needle but are able to generate and then retrieve it eventually. As shown in the Appendix Figure 10, Qwen2.5-14B-Instruct and Qwen2.5-72B-Instruct start by generating a refusal but then successfully retrieve the needle regardless. In contrast, the smaller Qwen2.5-7B-Instruct directly retrieves the needle without any refusal. This behavior explains the findings in Table 1 where Qwen2.5-14B-Instruct and Qwen2.5-72B-Instruct underperforming Qwen2.5-7B-Instruct when generation length is 32. Limiting the generation length only allows the larger Qwen models to generate the initial refusal and not the retrieved needle. Strangely, this refusal behavior is not observed when there are multiple inserted needles. Since this behavior is unique to the Qwen2.5 family and only for single Hybrid NIAH test, we hypothesize it is related to the Qwen’s chunked attention mechanism (An et al. 2024a) struggling to effectively focus on the singular inserted needle when generating the first token. When multiple needles are inserted, the syntactical similarity between the needles helps the larger Qwen models to focus on the context and they do not refuse to answer initially. We leave further investigation of this phenomenon for future work.
Conclusion
In this work, we investigate the role of parametric knowledge in LCLMs and uncover a critical trade-off between parametric recall and extrinsic retrieval. Our analysis reveals that enhancements achieved on extrinsic retrieval ability can inadvertently suppress the model’s use of its parametric knowledge, especially when the context is irrelevant or misleading. To evaluate this interplay, we introduce the novel Hybrid Needle-in-a-Haystack test that assesses a model’s ability to jointly integrate parametric and extrinsic knowledge. Experimental results demonstrate that even large-scale LCLMs struggle to effectively combine these two abilities and often fail to fully leverage their parametric knowledge during long-context generation. Our findings highlight the need for future model designs and evaluation protocols that consider both parametric and extrinsic knowledge sources, paving the way for more robust and context-aware LCLMs.
@@ Appendix A Hybrid Needle-in-a-Haystack Knowledge Dataset
All entities used in Hybrid NIAH tests are illustrated in Table  3, with the bolded entries representing the target entities used for generating the desired answers. To construct the random facts required for our Hybrid NIAH tests, we randomly select entities from all entities other than the current target entity.